OpenStudy (anonymous):

Solve the equation. -6p - 21 = 3p - 12 A. 1 B. 3 C. –3 D. –1 im thinking its 1 not really sure

OpenStudy (anonymous):

can't be 1

OpenStudy (anonymous):

\[-6p - 21 = 3p - 12\] if \(p=1\) you get \[-6-21=3-12\] which is not right

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

Hint: watch your negatives. :)

OpenStudy (anonymous):

thnx guys